OpenStudy (anonymous):

Determine if the integral converges or not and show steps : integral 1 to pi/2 tan(x) dx

OpenStudy (anonymous):

I really need help

OpenStudy (anonymous):

find the antiderivative of tan(x) first

OpenStudy (anonymous):

which is?

OpenStudy (anonymous):

(solve the integral without plugging limits first)

OpenStudy (anonymous):

can you work the thing out for me?

OpenStudy (anonymous):

tan(x) = sin(x)/cos(x) let cos(x) = u then du = -sin(x)dx so tan(x)dx = -du/u right ?

OpenStudy (anonymous):

yes ok continue

OpenStudy (anonymous):

so now integral of (-du/u) is simply -ln(u) = -ln(cos(x))

OpenStudy (anonymous):

agree ?

OpenStudy (anonymous):

so far yes now what

OpenStudy (anonymous):

it is in fact -ln|cos(x)|

OpenStudy (anonymous):

yes ok

OpenStudy (anonymous):

now we have to plug limits but doing so carefully .. by carefully i mean taking the limits

OpenStudy (anonymous):

ok so how

OpenStudy (anonymous):

as you can see lim [-ln|cos(x)| ] as x-> pi/2 is not converges right ?

OpenStudy (anonymous):

yes i see

OpenStudy (anonymous):

its divergent but at what number?

OpenStudy (anonymous):

then the whole integral wont converge

OpenStudy (anonymous):

so whats our final conclusion?

OpenStudy (anonymous):

not converges

OpenStudy (anonymous):

at what number? no number?

OpenStudy (anonymous):

give up.

OpenStudy (anonymous):

You cannot DIVERGE at a number - \[ \Huge \color{green} {\text{You can only diverge to infimity}} \pm \]

OpenStudy (anonymous):

I like this \[ \color{#F87431}{\Huge Expressionism\text{ of COLOR}}\]

OpenStudy (anonymous):

IT drains out the unneeded fatigue ....

OpenStudy (anonymous):

ITT: we Devolve to Inanity

OpenStudy (anonymous):

Did you know that the measure of finitely provable theorems in all true theorem is very LOW

OpenStudy (anonymous):

this integral diverges to 42/(6*7); equivalent to answer/(question)

OpenStudy (anonymous):

Read Godel and Church

OpenStudy (anonymous):

Madam ! My name is Church! Alonzo Church, but you can call me Alonzo... http://en.wikipedia.org/wiki/Alonzo_Church

OpenStudy (anonymous):

But this guy is the MOST INSPIRING and soul elevating of all Logicians http://en.wikipedia.org/wiki/Raymond_Smullyan \[ \LARGE \text{"The LAdy and the Tiger" by Raymond Smullyan}\]

OpenStudy (anonymous):

Elvis has left the building